# Bulk–boundary correspondence from hyper-invariant tensor networks

### v1

@@ Abstract
We introduce a hyper-invariant tensor network designed to faithfully simulate the AdS/CFT correspondence, akin to the multi-scale entanglement renormalization ansatz (MERA). The proposed construction integrates bulk indices within the network architecture to uphold the key features of the HaPPY code, including complementary recovery. This hyper-invariant framework accurately reproduces the boundary conformal field theory’s (CFT) two- and three-point correlation functions, while considering the image of any bulk operator. Furthermore, we provide an explicit methodology for calculating the correlation functions in an efficient manner. Our findings highlight the physical aspects of the relation between bulk and boundary within the tensor network models, contributing to the understanding and simulation of holographic principles in quantum information.
1 Introduction
@@
Holographic quantum error-correction codes are a modern approach to study the deep connection between the notion of AdS/CFT correspondence, introduced in Maldacena1998 ; Witten1998 , and quantum information. In recent years, numerous significant ideas have emerged, for a recent review see Jahn2021 . Among the most important of these ideas are HaPPY codes HaPPy_code , which rely on the mapping the bulk code space to the boundary by the isometry $\mathcal{V}:\mathcal{H}_{bulk}\xrightarrow{}\mathcal{H}_{boundary}$. The map $\mathcal{V}$ is realized through a network of perfect tensors perfect_tensors , which can be equivalently identified with absolutely maximally entangled (AME) states Goyeneche:2015fda ; quantum_perfect_tensors .
@@
Because HaPPY codes are placed on the hyperbolic tiling of the Poincare disk and map bulk degrees of freedom to the boundary ones, they were seen as the first connection between holography and quantum error correction. Moreover, it was proven that HaPPY codes preserve the area law ECP10 , as the Ryu–Takayanagi formula for the entropy dependence between bulk and boundary is satisfied Ryu2006 ; HaPPy_code .
Unfortunately, this model is too ideal to simulate any physically interesting scenario, because due to the perfect tensor symmetries, any two-point correlations on the boundary become trivial ChunJun2022 , thus do not depend on the distance between the points analyzed.
@@
From a physical standpoint, the main issue is that perfect tensor networks are too symmetric to encompass spatial directions and to reproduce realistic correlation functions between spatially separated points at the boundary. In contrast entirely different tensor networks, without bulk indices, called MERA, can indeed simulate the conformal field theory by producing correct two- and three-point correlation functions Pfiefer2009 . Moreover, within this approach, one can obtain the central charge of the Virasoro algebra Holzhey1994 ; Pfiefer2009 , using the von Neumann entropy in the continuum limit.
@@
Motivated by MERA networks, in 2017, Evenbly introduced the concept of hyper-invariant tensor networks (HTN) Evenbly2017 . This pioneering work studied the entanglement properties of HTNs and introduced a suitable discrete counterpart of scaling, thereby showing a connection with CFT. Using this technique various quantities such as the central charge $c$ and scaling dimensions of primary fields were computed ChunJun2022 ; Jahn2021 . An alternative approach Jahn2022 , aims to create CFT directly on the tensor network. This method extends the continuum notion of CFT to a discrete version, called Quasi-CFT (qCFT). Following this direction, the work Steinberg2022 proven that the construction of MERA-type tensor networks on hyperbolic tiling is at least a close approximation of continuum CFT, which is considered an example of qCFT Jahn2022 . However, to the best of our knowledge, no previous construction captured bulk-boundary mapping and presented a rigorous discussion of correlations on the boundary.
@@
The aim of this work is to introduce a tunable modification of the HaPPY code, which decreases the degree of its symmetry, but leads to realistic correlation functions. The proposed model connects them with nodes of hyper-invariant tenor networks Steinberg2023 , constructed from dual unitary matrices AWGG16 ; BKP19 ; BRRL24 . Such a technique allows us to combine the benefits of a HaPPy code – bulk to boundary mapping, with the hyper-invariant ones – desired forms of correlation functions. We note that the underlying utilized property of hyper-invariant tensor network nodes is planar $k$-uniformity planar_k_uniform_states .
@@
Within the presented approach, we managed to prove the desired behaviour of two- and three-point correlation functions in large network limits, while considering the image of any bulk operator. To derive these results we developed a methodology to conveniently and explicitly calculate desired correlation functions in an efficient manner. Due to the freedom in choosing the building blocks in our construction, we managed to numerically scan a wide range of scaling dimensions $\Delta$, demonstrating the potential to tailor our model to a specific value, corresponding to physically interesting fields.
@@
A HaPPY code HaPPy_code is determined by the geometry of the network and is applicable for any dimension $D$, provided a perfect tensor, with $n$ indices running from $1$ to $D$ does exist, which is equivalent to existence of AME$(n,D)$ states HESGW18 ; HW19b . It is often sufficient to take $D=2$ and work with qubits. For instance, the original code HaPPy_code is based on a perfect tensor with six indices each running from 1 to 2, which corresponds to an AME state of six qubits. However, choosing a larger local dimension $D$ significantly increases the space of possibilities. As the construction of hyper-invariant code Evenbly2017 employ at least pairs of qubits $d^{2}=4$, we present a suitable network also in the setting $D=d^{2}=4$. However, any other local dimension $d$ is suitable as well.
@@
2 Tensor networks construction
In this Section, we present a new construction of a hyper-invariant tensor network realizing the isometry between the Bulk and Boundary Hilbert spaces. Similarly, as well-established HaPPy codes HaPPy_code and hyper-invariant tensor networks Evenbly2017 , we place our structure on a tessellation of Poincare disc, which corresponds to a constant time slice of Anti-de-Sitter (AdS) space of $2+1$ dimension. We choose standard $\{5,4\}$ tiling of regular pentagons, four of which meet at each vertex, see Fig. 3 (a). This frame is filled, layer by layer, with tensors with one bulk index pointing "up" and five boundary indices, which contract with neighbouring tensors. The free boundary indices in the last filled layer correspond to local subspaces of the boundary Hilbert space.
@@ 2 Tensor networks construction
In order to reconstruct the physical properties of AdS, we impose additional "rotation symmetry" conditions for a perfect tensor – the cyclic change of tensor indices, except the first, bulk one, should remain invariant. i.e. $T_{ijklmn}=T_{injklm}$ – see Fig. 1(a). The symmetry of translations corresponds to the requirement for all nodes in the tensor network to be identical. As an exemplar tensor of this form, we chose a perfect tensor of $6$ quarts ($D=d^{2}=4$), corresponding to AME$(6,4)$ state, constructed with three orthogonal Latin hypercubes of order four Goyeneche:2015fda , since its local preprocessing by permutation matrices is sufficient to obtain desired form.
The second element of our construction "frame" $F$ is inspired by a node of a hyper-invariant tensor network. Its key property is being planar $2$-uniform planar_k_uniform_states : its contraction with its hermitian conjugate on $3$ neighbour pair of indices reduces it to identities on remaining pairs. Since the exact definition of a hyper-invariant tensor network is quite complex, and not essential in further discussion we presented in the appendix B. We propose a construction of such tensor in the spirit of Evenbly2017 with five copies of the dual-unitary matrix, see Figure 1 (b) and (c), which is why we call it hyper-invariant. To guarantee desired reductions of hyper-invariant tensor frame we indeed need only two simple identities, corresponding to two types of orthogonalities presented in Figure 2. Matrix with such properties can be realised on pair of qubits – ququart ($d^{2}=4$), see example in Sec. 4.
@@
Figure 1: (a) Diagrammatic representation of a perfect tensor obtained from 3 orthogonal Latin hypercubes of order four – see Fig. 9. All but the first index point in the "boundary" direction. Appropriate permutation matrices applied on outgoing indices make this tensor "rotationally invariant" with respect to rotation by $72^{\text{o}}$ degrees. In this and future graphics each line represents one qubit, the circles correspond to qubits pointing inwards. (b) Hyper-invariant tensor "frame" constructed from five $2$-qubit unitaries. (c) Desired contraction rule of hyper-invariant frame planar_k_uniform_states , pale colour corresponds to conjugation. Contraction of this tensor with its hermitian conjugate on $3$ neighbour pair of indices reduces it to identities.
@@
Figure 2: Necessary properties for blue dual-unitaries. The pale colour corresponds to hermitian conjugation. Each line corresponds to one qubit index.
The final step in our model is to combine perfect tensor $T$ with hyper-invariant frame $F$ into one node $N$ by entangling them on outgoing indices with large unitaries, as presented in Figure 3 (b). This step is crucial because otherwise entire tensor network would split up into two uninteresting networks, the first consisting of perfect tensors would be just a HaPPy code, and the second, hyper-invariant one, would have no connection to the bulk. In the discussed example these unitaries act on pairs of ququarts, $d^{4}=16$.
@@
We note, that our network construction, consisting of the perfect tensor $T$ together with hyper-invariant frames $F$ as nodes and pairs of unitaries on edges, would establish a relaxed version of the Evenbly code Steinberg:2024ack , see Appendix B. Furthermore, without perfect tensors $T$ it is a direct realization of a hyper-invariant tensor network Evenbly2017 , presented in Appendix B as well.
Constructed nodes have indeed desired reduction rules: The result of contracting the tensor with its conjugate on three neighbouring boundary indices is an identity on all remaining pairs of indices. This allows us to build a hyper-invariant tensor network, with appropriate reduction rules, but with possible non-vanishing 2-point correlation functions on the boundary.
Throughout the paper, we consider a tensor network built in a vertex inflation manner. In the zeroth step, the network consists of only one node. In each consecutive step new tensor is added to the free boundary index of the previous layer and then an additional tensors are added to close each vertex with already three tensors, as presented in Figure 3 (a).
Note that while constructing a tensor network in such a regular way, each new tensor, when added, has at least $3$ neighbouring non-contracted boundary indices, which is more than necessary to map the state from bulk and $2$ contracted boundary indices. Thus the proposed construction is indeed an isometry from bulk into the boundary, by the same token as perfect tensors are. Using the same arguments one can see that we chose the simplest possible tiling suitable for our construction. Indeed, if the basic tile would be a square, instead of the pentagon, then the node which closes a loop around a vertex has only $2$ free boundary indices out of $4$ and the bulk one, thus it cannot be an isometric by itself, since it maps from the larger space into a smaller one. We note, that in our discussion we omit the normalization of a hyper-invariant tensor as an isometry, to later restore it, if necessary.
@@
Figure 3: (a) Example of a constructed tensor network with $2$ layers and the central node.
(b) Entire construction of hyper-invariant tensor. The green pentagon corresponds to the perfect tensor, blue bipartite matrices to the hyper-invariant frame and red unitaries entangle two elements in one tensor node. To clarify the figure we colored the indices on which dual-unitaries act in blue.
3 Properties of proposed construction
@@
Because conformal theories are naturally based in continuous geometry settings, it is not yet well understood, whether we can consistently speak about the Virasoro algebra in the discrete setting. Nevertheless, some attempts were made to construct the discrete theories, called qCFT Jahn2022_1 .
Fortunately, the work Holzhey1994 established a relation between central charge, geodesic length, and von Neumann entropy for any conformal field theory. This allows us to compute the central charge by the discrete version of conformal scaling – adding layers to the tensor network, and computing the difference between entropy. The discussion and application of this approach for a hyperbolic tilling of Poincare disc (corresponding to AdS (2+1) space) can be found in Evenbly2014 ; Evenbly2011 ; Jahn:2019nmz ; Jahn2022 . Using Eq. (23) from Central_charges_and_scaling we provide an upper bound for the central charge in our theory:
@@
We stress the inequality sign, which results from the extrapolation of entanglement between each neighbouring pair nodes to be maximal. This property was proven for the HaPPy codes HaPPy_code , however, in the case of hyper-invariant tensors, up to our knowledge, it is still a hypothesis. Furthermore, our numerical calculations suggest that the actual value of the central charge $c_{\{5,4\}_{v}}$ is close to the upper bound (1). A large value of the central charge provides an argument, that our model does not encompass the so-called minimal conformal field theories, for which $0<c\leq 1$ Ginsparg:1988ui .
We start the discussion of correlation functions by invoking a well-known fact, that conformal invariance of conformal field theories fixes the structure of two- and three-point correlation functions Ginsparg:1988ui . Let $\phi(x)$ be to quantum field acting on states in corresponding Hilbert space. Let $\ell_{ij}$ be a distance between two chosen points calculated in conformal theory – along the boundary of AdS. Thus, if the theory is conformal invariant, the correlation of field in points $x_{1}$ and $x_{2}$ is given by:
@@
where the proportionality constant can be set to $1$ by normalization of fields. The parameter $\Delta$ is called a scaling dimension for a field and characterises its transformation properties. The three-point correlations on the other hand are given by
$\langle\phi(x_{1})\phi(x_{2})\phi(x_{2})\rangle=\frac{C_{123}}{\ell_{12}^{\Delta}\,\,\,\ell_{23}^{\Delta}\,\,\,\ell_{13}^{\Delta}},$ (3)
@@
Moreover, we will focus on probing operators with zero traces. To justify this choice, let us consider a one-point correlation function. Note, that while taking trace one can, in the first step reduce all tensors with $3$ boundary indices, resulting in a new boundary to reduce in the consecutive step. Repeating this procedure, by the same token as for perfect tensors, one obtains a very simple form of a one-point correlation function
@@
3.1 Two point correlation function
For HaPPY codes, while calculating 2-point correlation function, one may reduce the entire tensor network, as shown in ChunJun2022 , and the two points of interest are separated. Thus if we sent $1$-point correlations to zero (6), the $2$-point correlations resulting from HaPPY codes is trivial and factorize,

### v2

# Bulk-boundary correspondence from Hyperinvariant tensor networks

@@ Abstract
We introduce a tensor network designed to faithfully simulate the AdS/CFT correspondence, akin to the multi-scale entanglement renormalization ansatz (MERA), following hyper-invariant tensor network. The proposed construction integrates bulk indices within the network architecture to uphold the key features of the HaPPY code, including complementary recovery. This framework accurately reproduces the boundary conformal field theory’s (CFT) two- and three-point correlation functions, while considering the image of any bulk operator. Furthermore, we provide an explicit methodology for calculating the correlation functions in an efficient manner. Our findings highlight the physical aspects of the relation between bulk and boundary within the tensor network models, contributing to the understanding and simulation of holographic principles in quantum information.
tensor networks, holography, bulk-boundary correspondence, correlations
I Introduction
@@
Holographic quantum error-correction codes are a modern approach to study the deep connection between the notion of AdS/CFT correspondence, introduced in [1, 2], and quantum information. In recent years, numerous significant ideas have emerged, for a recent review see [3]. Among the most important of these ideas are HaPPY codes [4], which rely on the mapping the bulk code space to the boundary by the isometry $\mathcal{V}:\mathcal{H}_{bulk}\xrightarrow{}\mathcal{H}_{boundary}$. The map $\mathcal{V}$ is realized through a network of perfect tensors [5], which can be equivalently identified with absolutely maximally entangled (AME) states [6, 7].
@@
Because HaPPY codes are placed on the hyperbolic tiling of the Poincare disk and map bulk degrees of freedom to the boundary ones, they were seen as the first connection between holography and quantum error correction. Moreover, it was proven that HaPPY codes preserve the area law [8], as the Ryu–Takayanagi formula for the entropy dependence between bulk and boundary is satisfied [9, 4]. Unfortunately, this model is too ideal to simulate any physically interesting scenario, because due to the perfect tensor symmetries, any two-point correlations on the boundary become trivial [10], thus do not depend on the distance between the points analyzed.
@@
From a physical standpoint, the main issue is that perfect tensor networks are too symmetric to encompass spatial directions and to reproduce realistic correlation functions between spatially separated points at the boundary. Some approaches to this problem were based on small "distortions" of perfect tensors [11, 12], in which the first order of two-point correlation’s expansion exhibits the desired behaviour. In contrast entirely different tensor networks, without bulk indices, called MERA, can indeed simulate the conformal field theory by producing correct two- and three-point correlation functions [13]. Moreover, within this approach, one can obtain the central charge of the Virasoro algebra [14, 13], using the von Neumann entropy in the continuum limit.
@@
Motivated by MERA networks, Evenbly introduced the concept of hyper-invariant tensor networks (HTN) [15]. This pioneering work from 2017 studied the entanglement properties of HTNs and introduced a suitable discrete counterpart of scaling, thereby showing a connection with CFT. Using this technique various quantities such as the central charge $c$ and scaling dimensions of primary fields were computed [10, 3]. An alternative approach [16], aims to create CFT directly on the tensor network. This method extends the continuum notion of CFT to a discrete version, called Quasi-CFT (qCFT). Following this direction, the work [17] proven that the construction of MERA-type tensor networks on hyperbolic tiling is at least a close approximation of continuum CFT, which is considered an example of qCFT [16]. However, to the best of our knowledge, no previous construction captured bulk-boundary mapping and presented a rigorous discussion of correlations on the boundary.
@@
The aim of this work is to introduce a tunable modification of the HaPPY code, which decreases the degree of its symmetry, but leads to realistic correlation functions. The proposed model connects them with nodes of hyper-invariant tenor networks [15, 17], constructed from dual unitary matrices [18, 19, 20]. Such a technique allows us to combine the benefits of a HaPPY code – bulk to boundary mapping, with the hyper-invariant ones – desired forms of correlation functions.
The proposed construction satisfies the definition of block-perfect tensors [21], which form a particular case of planar $k$-uniform tensors [22]. However, as the present work was motivated by hyper-invariant tenor networks and hyper-invariant codes [23], we prefer to stick to this nomenclature.
@@
Within the presented approach, we managed to prove the desired behavior of two- and three-point correlation functions in large network limits, while considering the image of any bulk operator. To derive these results we developed a methodology to conveniently and explicitly calculate desired correlation functions in an efficient manner. Due to the freedom in choosing the building blocks in our construction, we managed to numerically scan a wide range of scaling dimensions $\Delta$, demonstrating the potential to tailor our model to a specific value, corresponding to physically interesting fields.
@@
A HaPPY code [4] is determined by the geometry of the network and is applicable for any dimension $D$, provided a perfect tensor, with $n$ indices running from $1$ to $D$ does exist, which is equivalent to existence of AME$(n,D)$ states [24, 25]. It is often sufficient to take $D=2$ and work with qubits. For instance, the original code [4] is based on a perfect tensor with six indices each running from 1 to 2, which corresponds to an AME state of six qubits. However, choosing a larger local dimension $D$ significantly increases the space of possibilities. As the construction of hyper-invariant tensor network [15] employ at least pairs of qubits $d^{2}=4$, we present a suitable network also in the setting $D=d^{2}=4$. However, any other local dimension $d$ is suitable as well.
@@
II Tensor networks construction
In this Section, we present a new construction of a network realizing the isometry between the Bulk and Boundary Hilbert spaces. Similarly, as well-established HaPPY codes [4] and hyper-invariant tensor networks [15], we place our structure on a tessellation of Poincare disc, which corresponds to a constant time slice of Anti-de-Sitter (AdS) space of $2+1$ dimension. We choose standard $\{5,4\}$ tiling of regular pentagons, four of which meet at each vertex, see Fig. 3 (b). This frame is filled, layer by layer, with tensors with one bulk index pointing "up" and five boundary indices, which contract with neighboring tensors. The free boundary indices in the last filled layer correspond to local subspaces of the boundary Hilbert space.
@@ II Tensor networks construction
In order to reconstruct the physical properties of AdS space, we impose additional "rotation symmetry" conditions for a perfect tensor – tensor should remain invariant under cyclic change of tensor indices, except the first, bulk one, i.e. $T_{ijklmn}=T_{injklm}$ – see Fig. 1(a). The symmetry of translations corresponds to the requirement for all nodes in the tensor network to be identical. As an exemplar tensor of this form, we chose a perfect tensor of $6$ quarts ($D=d^{2}=4$), corresponding to AME$(6,4)$ state, constructed with three orthogonal Latin hypercubes of order four [6], since its local preprocessing by permutation matrices is sufficient to obtain desired form.
The second element of our construction "frame" $F_{abcde}$ is inspired by a node of a hyper-invariant tensor network. Its key property is being planar $2$-uniform [22]: its contraction with its hermitian conjugate on $3$ neighbor pair of indices reduces it to identities on remaining pairs. Since the exact definition of a hyper-invariant tensor network is quite complex, and not essential in further discussion we presented in the Appendix B. We propose a construction of such tensor in the spirit of [15] with five copies of the dual-unitary matrix, see Figure 1 (b) and (c), which is why we call it hyper-invariant. To guarantee desired reductions of hyper-invariant tensor frame we need only two simple identities, corresponding to two types of orthogonalities, satisfied by dual-unitary matrices, presented in Figure 2. Matrix with such properties can be realized on pair of qubits – ququart ($d^{2}=4$), see example in following sections.
@@
Figure 1: (a) Diagrammatic representation of a perfect tensor of six subsystems of dimension $D=4$ obtained from $3$ orthogonal Latin hypercubes of order four – see Fig. 11. The first index $i$, pointing in the "bulk direction", is denoted as a dot, while all the others indices, $j,\dots,n$, point in the "boundary" direction. Following example from the text, each line represents a single qubit, so a ququart index is represented by a double line. Appropriate permutation matrices applied on outgoing indices make this tensor ’rotationally invariant’ with respect to rotation of the pentagon by $52\pi/$ 5. (b) Hyper-invariant tensor "frame" constructed from five unitaries of order $D=4$ (blue rectangles). (c) Desired contraction rule of hyper-invariant frame [22], pale color corresponds to conjugation. Contraction of this tensor with its Hermitian conjugate on $3$ neighbor pair of indices reduces it to identities.
@@
Figure 2: Necessary properties for blue dual-unitaries. The pale colour corresponds to hermitian conjugation.
The final step in our model is to combine perfect tensor $T$ with hyper-invariant frame $F$ into one node $N$ by entangling them on outgoing indices with large unitaries, as presented in Figure 3 (b). This step is crucial because otherwise the entire tensor network would split up into two uninteresting networks, the first consisting of perfect tensors would be just a HaPPY code, and the second, hyper-invariant one, would have no connection to the bulk. In the discussed example these unitaries act on pairs of ququarts, $D^{2}=d^{4}=16$.
@@
Constructed nodes have profitable reduction rules: The result of contracting the tensor with its conjugate on three neighboring boundary indices is an identity on all remaining pairs of indices. This allows us to build a hyper-invariant tensor network, with appropriate reduction rules, but with possibly non-vanishing correlation functions on the boundary.
Throughout the paper, we consider a tensor network built in a vertex inflation manner. In the zeroth step, the network consists of only one node. In each consecutive step new tensor is added to the free boundary index of the previous layer and then an additional tensors are added to close each vertex with already three tensors, as presented in Figure 3 (b). While constructing a tensor network in such a regular way, each new tensor, when added, has at least $3$ neighboring non-contracted boundary indices, which is more than necessary to serve as isometry from bulk and $2$ contracted boundary indices. Thus the proposed construction is indeed an isometry from bulk into the boundary, by the same token as perfect tensors are. Using the same arguments one can see that we chose the simplest possible tiling suitable for our construction. If the basic tile would be a square, instead of the pentagon, then the node which closes a loop around a vertex has only $2$ free boundary indices out of $4$ and the bulk one, thus it cannot be an isometric by itself, since it maps from the larger space into a smaller one. In our discussion we omit the annoying normalization of a hyper-invariant tensor as an isometry, to later restore it, if necessary.
@@
Figure 3: (a) Entire construction of a hyper-invariant tensor. Green pentagon corresponds to a perfect tensor with six indices running from 1 to $D=4$, blue rectangle denote dual-unitary matrices of order $D$, forming a hyper-invariant frame, while red unitaries of size $D^{2}=16$ entangle two elements in a single tensor node. Each line represents a single qubit, blue lines represent indices on which dual-unitaries act. The bulk index, pointing "inward", is denoted as a dot inside a green pentagon. (b) Example of a constructed tensor network with $2$ layers and the central node.
Constructed network based on a perfect tensor $T$, a with hyper-invariant frames $F$ as nodes and pairs of unitaries on edges presented in Fig. 4 leads to a relaxed version of an Evenbly code [26] – see Appendix B. An analogous construction without perfect tensors forms a direct realization of a hyper-invariant tensor network [15], presented in Appendix B.
Figure 4: Connection of $2$ nodes from the tensor network. A pair of unitaries (rectangles in red) from neighboring nodes can be considered as an edge tensor (red circle) and a perfect tensor with a frame as a node tensor (blue circle) from hyper-invariant code.
The perspective of hyper-invariant codes naturally indicates gauge freedom in a tensor network constructed from presented nodes. Denoting entangling unitary gate (red rectangle) as $U$, we obtain the product $UU^{\top}$ at each edge of the tensor network – see Fig 4. Replacing $U$ by its product with any orthogonal matrix $U\to UQ$, does not change the behavior of the network as $(UQ)(UQ)^{\top}=UQQ^{\top}U^{\top}=UU^{\top}$. One needs only o take into account the uncontracted boundary indices on which the action of $Q$ does not disappear.
III Properties of construction proposed
@@
Because conformal theories are naturally based in continuous geometry settings, it is not yet well understood, whether we can consistently speak about the Virasoro algebra in the discrete setting. Nevertheless, some attempts were made to construct the discrete theories, called qCFT [27]. Fortunately, the work [14] established a relation between central charge, geodesic length, and von Neumann entropy for any conformal field theory. This allows us to compute the central charge by the discrete version of conformal scaling – adding layers to the tensor network, and computing the difference between entropy. The discussion and application of this approach for a hyperbolic tilling of Poincare disc (corresponding to AdS (2+1) space) can be found in [28, 29, 30, 16]. Using Eq. (23) from [31] we provide an upper bound for the central charge in our theory:
@@
We stress the inequality sign, which results from the extrapolation of entanglement between each neighboring pair nodes to be maximal. This property was proven for the HaPPY codes [4], however, in the case of hyper-invariant tensors, up to our knowledge, it is still a hypothesis. Furthermore, our numerical calculations suggest that the actual value of the central charge $c_{\{5,4\}_{v}}$ is close to the upper bound (1). A large value of the central charge provides an argument, that our model does not encompass the so-called minimal conformal field theories, for which $0<c\leq 1$ [32].
We start the discussion of correlation functions by invoking a well-known fact, that conformal invariance of conformal field theories fixes the structure of two- and three-point correlation functions [32]. Let $\phi(x)$ be to quantum field acting on states in corresponding Hilbert space. Let $\ell_{ij}$ be a distance between two chosen points calculated in conformal theory – along the boundary of AdS space. Thus, if the theory is conformal invariant, the correlation of field in points $x_{1}$ and $x_{2}$ is given by:
@@
where the proportionality constant can be set to $1$ by normalization of fields. The parameter $\Delta$ is called a scaling dimension for a field and characterizes its transformation properties. The three-point correlations on the other hand are given by
$\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle=\frac{C_{123}}{\ell_{12}^{\Delta}\,\,\,\ell_{23}^{\Delta}\,\,\,\ell_{13}^{\Delta}},$ (3)
@@
Moreover, we will focus on probing operators with zero traces. To justify this choice, let us consider a one-point correlation function. Note, that while taking trace one can, in the first step reduce all tensors with $3$ boundary indices, resulting in a new boundary to reduce in the consecutive step. Repeating this procedure, by the same token as for perfect tensors, one obtains a very simple form of a one-point correlation function,
@@
III.1 Two point correlation function
For HaPPY codes, while calculating 2-point correlation function, one may reduce the entire tensor network, as shown in [10], and the two points of interest are separated. Thus if we sent $1$-point correlations to zero (6), the $2$-point correlations resulting from HaPPY codes is trivial and factorize,